OpenStudy (anonymous):

prove or disprove , for any prime p>2 there exist an integer n such that :- n(n+1)

OpenStudy (kainui):

Divide both sides by (n+1) and you get n>n+2 so all primes? XD

OpenStudy (anonymous):

i'll modify the question

OpenStudy (anonymous):

what do u think of it nw :P

OpenStudy (kainui):

\[ n(n+1)<p<(n+1)(n+2)\]Ok now divide by (n+1) again and we have \[ n<p<n+2\] so if we let p=n+1 we have \[n<n+1<n+2\] which are all consecutive numbers. So to be true, we just pick n=p-1 and this will always work I think.

OpenStudy (anonymous):

:P yes was gonna troll by letting u guys trying to find counter example xD

OpenStudy (anonymous):

but since u got it fun is over :P

OpenStudy (kainui):

hahaha XD I was unsure cause it seemed easier than it should have been I was trying to find the counter example at first haha.

OpenStudy (anonymous):

the inverse though is not true :- " for any integer n there exist p prime , such that n(n+1)<p<(n+1)(n+2) "

OpenStudy (anonymous):

now try to disprove this :P

OpenStudy (anonymous):

:D just kidding lol

OpenStudy (kainui):

lol XD

OpenStudy (anonymous):

ikr :P

OpenStudy (anonymous):

so reached what as counter example?

OpenStudy (kainui):

lol no idea

OpenStudy (kainui):

I think I need to find a gap of 2n+2 composite numbers to find n.

OpenStudy (anonymous):

the thing is as n become bigger the gap btw n(n+1) and (n+1)(n+2) become bigger :D so sometimes we even have many primes btw think of it , i'll go and study :P good night :D

OpenStudy (kainui):

Haha yeah it's a big problem. I'll see what I can do. =D

ganeshie8 (ganeshie8):

*